Write the equation of the line that satisfies the given conditions. Express the final equation in standard form.
Contains the point (-7, 4) and is parallel to the line x − 4y = 6

Answer by Alan3354(69443)

Write the equation of the line that satisfies the given conditions. Express the final equation in standard form.
Contains the point (-7, 4) and is parallel to the line x − 4y = 6
--------
Find the slope of the line, m.
---
Parallel lines have the same slope.
Use y-y1 = m*(x-x1) where (x1,y1) is the point.